### v1

@@ 3 Spectral types and velocity determination
We performed an analysis to determine the spectral type and line-of-sight (l.o.s.) velocity of our sources. For this purpose, we used, respectively, the spectral fitting codes ULySS (Koleva et al. 2009) and spexxy (Husser 2012).
@@ 3.1 Spectral classification with ULySS
The spectral type classification was carried out with the same approach as in Roth et al. (2018); we refer the reader to this article for full details. A linear combination of up to 10 templates from the MIUSCAT stellar spectral library555The MIUSCAT stellar spectral library fills in the gap between the MILES (Sánchez-Blázquez et al. 2006) and the near-IR CaT library of Cenarro et al. (2001), using the Indo-US spectral library of Valdes et al. (2004); it also extends the wavelength coverage slightly blue-ward of MILES and red-ward of the CaT library. MIUSCAT has been recently replaced by E-MILES (Vazdekis et al. 2016). (Vazdekis et al. 2012) was fitted to the observed spectra using the ULySS666http://ulyss.univ-lyon1.fr code (Koleva et al. 2009). We note that ULySS is capable of identifying unresolved blended stars because the spectrum is fitted with a linear combination of non-linear components, multiplied by a polynomial continuum. In our case, no blends are found, not even in the central field. This is mostly expected, given the density of sources visible in Fig. 2, which was constructed from the MUSE data cubes.
@@
We stored the information of the three spectral types having the largest weights in the best fit. Occasionally, the fit returns implausible mixes of spectral types (e.g., mixing young early-types with evolved late-types). The apparent magnitude obtained by shifting the absolute magnitude in R-band of the stars in the MIUSCAT library to the distance of IC 1613 (distance modulus , from Bernard et al. 2010), provided a way to weed out these clearly erroneous fits when they would lead to apparent magnitudes well below the detection limit of the MUSE data. However, we also found that good-quality fits to cool MS stars allowed us to confidently identify Milky Way foreground stars that would otherwise have been difficult to reject from photometry alone.
@@
The ULySS spectral analysis was carried out on all those sources having a S/N and a clean photometry (see Sect. 2.1). Main outputs of the code were the best fitting templates, together with the associated spectral parameters (, log(g) and [Fe/H]) and l.o.s. velocity. We verified by eye the ULySS outcomes, assigning several quality flags to each analysed object. We evaluated the quality of the input spectrum (QSP), the quality of the spectral fitting (QFT), the plausibility of the output l.o.s. velocity (PVR) and the plausibility of the spectral type classification (PCL); flags were reported as integer numbers ranging from one to four, with the lower value meaning a useless measure, and the higher value implying an excellent fit. In this work, we used ULySS exclusively for spectral classification purposes, while we left the l.o.s. velocity determination part to the spexxy code. This was mainly because spexxy performed better than ULySS in the velocity determination task, especially in the low S/N regime (i.e., ), as we verified in Appendix B.1 and B.2.
@@
In Table 2, we report results from the spectral classification obtained for sources marked with a PCL flag between three and four (which implies an accuracy on the effective temperatures better than 500 K). The visual inspection during the spectral classification also allowed us to identify other contaminants that were not removed during the previous photometric selection (Sect. 2.1): they were mainly background galaxies (comprising high-z emitters) and low S/N targets highly polluted by diffuse ionised gas emission lines. We also visually identified carbon and B emission line stars that were missed by ULySS due to the lack of adequate stellar templates, in order to study them separately (see Sect. 3.3). We note that we define as bona-fide B emission-line stars (noted as Be hereafter) those sources whose spectral distribution is compatible with a star having K, but which show H and (occasionally) H in strong emission (e.g., Porter & Rivinius 2003; Rivinius et al. 2013). The other seven stars showed only H in weak emission, but due to their low S/N their classifications were too uncertain and they were therefore discarded.
Overall, we obtained a reliable spectral classification for 808 stars888The spectroscopic for the red giant stars were found to be in fair agreement with the empirical relation from Alonso et al. (1999), which is valid in the colour range and mostly metallicity independent.. Looking at Table 2, the large majority of sources are K-type giants, with an almost even representation of the other types, except for the M and O stars being a minority. In particular, among the M stars, only one was found to be a giant, while the other four were classified as MW-foreground dwarfs. We also report the presence of two other MW-contaminants classified as K V and G V stars, all confirmed in the l.o.s. velocity determination step. We have therefore excluded these MW-foreground stars from the counts in Table 2 and the following analysis. Fig. 4 shows the position of the classified stars on the colour-magnitude diagram, colour-coded according to their , visually confirming the general goodness of the spectral classification.
@@
On the other hand, O stars are typically not uniformly distributed in space. They tend to form and evolve in OB associations. Our observations cover some of the OB associations identified by Garcia et al. (2009). These associations are low density groups of young stars with significant internal extinction, making the identification of O types more challenging. In Fig. 5 we show the spatial distribution of the OB associations reported by Garcia et al. (2009), together with that of our identified OB stars with K and also the Be stars. We see that they generally tend to be found where the OB associations are. Their spatial distribution also coincides with the ionised shells shown in Fig. 2, to which they are probably physically associated (Borissova et al. 2004; Garcia et al. 2009). At this stage, these evidences at least confirm the goodness of our initial classification of many of them as hot stars.
@@
3.2 Velocity determination with spexxy
@@
The determination of l.o.s. velocities was carried out using the spexxy999https://github.com/thusser/spexxy code (v. 2.5, Husser 2012). The routine performs a full spectral fitting to the observed spectra using interpolated spectral templates generated from the PHOENIX library of high-resolution synthetic spectra101010http://phoenix.astro.physik.uni-goettingen.de/ (Husser et al. 2013). The library covers a wide wavelength range, going from 500 Å to 5.5 m, and stellar parameters: , , and . However, the PHOENIX models do not treat radiative transfer in the atmospheres of massive stars affected by strong stellar winds, so the library is of little use in the recovery of stellar parameters for K. We refer to Appendix B.1 for a detailed explanation of the spectral fitting steps and to B.2 for a comparison of its performance with the ULySS results presented above.
@@
We ran spexxy on all targets having S/N (in at least one of the S/N indicators) and clean photometry (see Sect. 2.1). We also excluded all those targets that in the spectral classification step (see Sect. 3.1) were identified as contaminated by ionised gas, background galaxies, and Be stars (which we analysed separately). These steps reduced our sample to 1745 input sources. We were thus able to recover their l.o.s. velocity, along with estimates of their spectral parameters and associated errors. We note that only for of stars spexxy failed to perform a successful fitting, mostly due to their low S/N.
@@
The goodness of the recovered parameters was quantified in a series of tests, whose details are given in Appendix B.1. Here we report the essentials. We verified by performing several tests on mock spectra that the l.o.s. velocity errors are well estimated down to S/N for red giant stars (S/N for MS stars), and the velocity values are recovered without significant offsets. On the other hand, the recovery of the spectral parameters was more limited, with in particular being generally well constrained for red giants, but underestimated for the hottest MS stars due to the grid limitation of our spectral models. We emphasise again that our aim was to use spexxy mainly for the determination of l.o.s. velocities.
@@
Since our main goal is to analyse not only the global kinematic properties of IC 1613, but also how they change as a function of stellar types, we decided to keep the sample obtained during the spectral classification step. From this sample we further removed those targets with a velocity error km s, which meant excluding of sources with the least reliable velocity values (mostly low S/N MS stars). We also excluded the six targets previously classified as MW contaminants, confirming that they are foreground sources based on their l.o.s. velocities, which were consistent with 0 km s. Performing a cross-correlation with the third data release of the Gaia catalogue (Gaia Collaboration et al. 2023), we further confirmed their foreground nature based on their proper motion values, significantly different from 0 km s in this case. With these conditions applied, our sample reduced to 727 sources to which we subsequently added the 24 Be stars whose velocity determination we describe in the next section. The distribution of l.o.s. velocity measurements for our sample is shown in Fig. 6, with histograms divided by pointing and stellar type. We note that the median velocity error was of km s, while the median S/N resulted around 10.
3.3 Velocity determination for carbon and Be stars
@@
The l.o.s. velocities of the C stars identified by spectral classification (see Sect. 3.1) were calculated using spexxy as for the main sample. However, the code was not always able to match the complex spectral features of these stars. Therefore, we double-checked the spexxy velocity determination with ULySS using stellar templates from the X-Shooter library of carbon stars (Gonneau et al. 2016). We found a general agreement within the errors, confirming the goodness of the spexxy velocities. The median velocity error of the 14 C stars analysed (6 in both F1 and F2, and 2 in F3) was km s for a median S/N.
@@
On the other hand, the velocity determination for the Be stars was done separately, since spexxy failed in this task because such stars are not represented by the PHOENIX model atmospheres. We thus performed a cross-correlation with custom-made template spectra using the function correlation of the python package specutils, part of the Astropy project (Astropy Collaboration et al. 2022). Templates were generated as ad-hoc proxies using the spectra of two B-type stars (classified as B9III and B3III) from the MIUSCAT library (Vazdekis et al. 2012), to which we added two Gaussian profiles reproducing the strength and width of the H and H emission lines as detected in our sample. Since the spectra of our Be-star candidates in general did not show any split line profiles (except for two objects), the two-component Gaussian approximation seemed to be a good enough approach for the purpose of l.o.s. velocity measurements. We cross-correlated the observed spectra with the templates and assigned them the average of the derived velocities. The errors were instead assigned by a Monte Carlo process. For each template we generated mock spectra at different S/N (calculated around 5500 Å) values . For each S/N, we cross-correlated 250 mock spectra with the corresponding noise-free template and considered the median absolute deviation of the velocity distribution as the velocity error representative of that bin. The error distribution as a function of S/N was fitted by an exponential profile for each template and we used the mean profile to assign velocity errors to the observed spectra according to their S/N. For the 24 Be stars analysed (14, 6 and 4 in fields F1, F2 and F3, respectively), we recovered a median km s for a median S/N.
@@ 4 Internal kinematics
| Sample | N | N | Model | k | Bayes factor | ||
| (km s) | (km s) | (km s arcmin) | ln() | ||||
| Main | 751 | 746 | Linear rotation (HI) | 4.5 | |||
| Linear rotation (optical) | 3.5 | ||||||
@@
We investigate the basic internal kinematic properties of the stellar component of IC 1613, such as its systemic l.o.s. velocity, its velocity dispersion and the possible presence of velocity gradients, indicating rotation of the stellar component121212On the angular scales covered by the MUSE sample, we find the perspective rotation due to the relative motion between the galaxy and the Sun to play a negligible role with up to km s (following Eq. 6 of van de Ven et al. 2006 and assuming the systemic proper motion from Battaglia et al. 2022). Velocity gradients induced by tidal disturbances from the MW or M31 are also highly unlikely..
@@
To proceed with the kinematic analysis, we first need to identify the possible contaminants left in our sample. We had already removed many of them during the spectral classification process, with the rest expected to be faint foreground stars. We used then the method outlined in Taibi et al. (2020), which allows a Bayesian kinematic analysis while assigning membership probabilities to the individual targets. This approach is based on the expectation maximization technique presented in Walker et al. (2009).
@@
The values depend on the l.o.s. velocities of the individual targets, but also on the prior information based on their radial distances from the galaxy’s centre and on the expected l.o.s. velocity distribution of the possible contaminants. The spatial prior takes into account that the probability of membership is higher, the closer to the galaxy’s centre. In this case, we simply required that our targets should follow a monotonically decreasing radial density profile applying an isotonic regression model, with no assumptions about its functional form131313https://scikit-learn.org/stable/modules/isotonic.html.
We used the Besançon model of Galactic foreground stars (Robin et al. 2003) to obtain the expected l.o.s. velocity distribution of the contaminants. We generated a catalogue along IC 1613’s direction, over an area up to its half-light radius, covering the range of colours and magnitudes of our targets. This resulted in a velocity distribution that was well sampled and representative of the spatial area around the MUSE pointings. We approximated it as the sum of two Gaussian profiles (with means km s, km s, and standard deviations km s, km s, with an amplitude ratio of ). According to this model, considering the small area covered by the MUSE pointings and the spanned range of magnitudes and colours, we expect to find, in total, around five foreground contaminants in the velocity range between km s and 200 km s. If we restrict to the velocity range covered by the input targets (i.e. between km s and km s, see Fig. 6), we would expect only one contaminant. We recall that already during the spectral classification step (see Sect. 3.1) we found and removed six likely contaminants, all being cold MS stars with velocities around .
@@
We carry out a Bayesian analysis to explore and compare different kinematic models, considering one where the internal kinematics of IC 1613’s stellar component can be purely described by random motions, and another model which also contains a rotational term. For the rotational term we used the form: , where k is the velocity gradient, are the angular distance from the galaxy’s centre (i.e. , where () are the tangent plane’s coordinates) and position angle (P. A.) of the i-target, and is the P. A. of the kinematic major axis (both P. A. measured from North to East).
@@
We note that, unless otherwise stated, for the stellar component we quote the observed rotational velocity and not its intrinsic value , as the latter requires knowledge of the inclination angle i of the angular momentum vector with respect to the observer (i.e., ). The stellar component of Local Group dwarf galaxies is not in a thin disc, therefore its inclination cannot be determined without knowledge of the intrinsic 3D shape. In those systems that contain a clearly rotating HI component, as IC 1613, one can at least assume the inclination determined from the neutral gas component, under the assumptions that stars and gas share the same inclination. This is what we will do when quoting values for or when deriving estimates for the circular velocity. However, we note the possibility that IC 1613 may be seen close to face-on (Read et al. 2016), therefore even the HI value has its limitations.

### v2

@@ 3 Spectral types and velocity determination
We performed an analysis to determine the spectral type and line-of-sight (l.o.s.) velocity of our sources. For this purpose, we used, respectively, the spectral fitting codes ULySS (Koleva et al. 2009)555http://ulyss.univ-lyon1.fr and spexxy (Husser 2012). The reason for this dual approach is motivated as follows.
Crowded field IFS can be considered an accomplished technique for globular clusters (GCs; Giesers et al. 2018, 2019; Latour et al. 2019; Kamann et al. 2020; Saracino et al. 2022). However, for galactic systems at larger distances, involving young stellar populations and ongoing star formation, the method is still at a stage of exploration. Unlike GCs that are dominated by old stars, complications related to hot stars, emission line stars, carbon stars, but also to unresolved star clusters and the presence of strong emission lines from H II regions, make it difficult at present to blindly use automated fitting tools such as spexxy for nearby galaxies. The scheme of visually assisted spectral fitting described below was first developed in a pilot study of the disk galaxy NGC 300 at a distance of 1.88 Mpc (Roth et al. 2018, hereafter R18).
@@ 3.1 Spectral classification with ULySS
3.1.1 General method and classification results
Following R18, we used ULySS to fit the observed spectra to a linear combination of 10 templates with different weights. Templates were extracted from the MIUSCAT empirical library666The MIUSCAT stellar spectral library fills in the gap between the MILES (Sánchez-Blázquez et al. 2006) and the near-IR CaT library of Cenarro et al. (2001), using the Indo-US spectral library of Valdes et al. (2004); it also extends the wavelength coverage slightly blue-ward of MILES and red-ward of the CaT library. MIUSCAT has been recently replaced by E-MILES (Vazdekis et al. 2016)., rather than from a model grid. We note that we supplemented the MIUSCAT library with 19 spectra of carbon stars from the X-Shooter library (Chen et al. 2014)777http://xsl.astro.unistra.fr/page\_dr1\_all.html. We also added a set of mock spectra of Be stars composed ad hoc from library B-star spectra, combining two-component Gaussians at Hα and Hβ wavelengths with different equivalent widths. While these modifications did not aim to derive significant stellar parameters, they proved useful in identifying candidate stars that are not included in the MIUSCAT library for their later visual confirmation and velocity determination (see Sect. 3.2.2).
@@ 3.1.1 General method and classification results
The linear combination fitting mode of ULySS has been employed in R18 with the experience that unresolved blends from different spectral types are flagged to prompt an inspection. However, we found that the mild crowding in IC 1613 did not present us with such cases, unlike the more distant galaxy NGC 300. Occasionally though, the fit returned implausible mixes of spectral types (e.g., mixing young early-types with evolved late-types) for low S/N stars. The apparent magnitude obtained by shifting the absolute magnitude in R-band of the stars in the MIUSCAT library to the distance of IC 1613 (distance modulus , from Bernard et al. 2010), provided a way to weed out these clearly erroneous fits when they would lead to apparent magnitudes well below the detection limit of the MUSE data. We also found that good-quality fits to cool MS stars allowed us to confidently identify Milky Way foreground stars that would otherwise have been difficult to reject from photometry alone.
@@
The ULySS spectral analysis was carried out on all those sources having a S/N and a clean photometry (see Sect. 2.2). Main outputs of the code were the best fitting templates, together with the associated spectral parameters (, log(g) and [Fe/H]) and l.o.s. velocity. We verified by eye the ULySS outcomes, storing the information of the most probable template chosen among those with the largest weights. We also assigned several quality flags to each analysed object, evaluating the quality of the input spectrum (QSP), the quality of the spectral fitting (QFT), the plausibility of the output l.o.s. velocity (PVR) and the plausibility of the spectral type classification (PCL). Flags were reported as integer numbers ranging from one to four, with the lower value meaning a useless measure, and the higher value implying an excellent fit. We refer the reader to R18 for further details.
In this work, we used ULySS exclusively for spectral classification purposes, while we left the l.o.s. velocity determination part to the spexxy code. This was mainly because spexxy performed better than ULySS in the velocity determination task, especially in the low S/N regime (i.e., ), as we verified in Appendix B.1 and B.2.
@@
In Table 2, we report results from the spectral classification obtained for sources marked with a PCL flag between three and four (which implies an accuracy on the effective temperatures better than 500 K). The visual inspection during the spectral classification also allowed us to identify other contaminants that were not removed during the previous photometric selection (Sect. 2.2): they were mainly background galaxies (comprising high-z emitters) and low S/N targets highly polluted by diffuse ionised gas emission lines. We also visually identified carbon and B emission line stars that were missed by ULySS due to the lack of adequate stellar templates, in order to study them separately (see Sect. 3.2.2). We note that we define as bona-fide B emission-line stars (noted as Be hereafter) those sources whose spectral distribution is compatible with a star having K, but which show Hα and (occasionally) Hβ in strong emission (e.g., Porter & Rivinius 2003; Rivinius et al. 2013). The other seven stars showed only Hα in weak emission, but due to their low S/N their classifications were too uncertain and they were therefore discarded.
Overall, we obtained a reliable spectral classification for 808 stars999The spectroscopic for the red giant stars were found to be in fair agreement with the empirical relation from Alonso et al. (1999), which is valid in the colour range and mostly metallicity independent.. Looking at Table 2, the large majority of sources are K-type giants, with an almost even representation of the other types, except for the M and O stars being a minority. In particular, among the M stars, only one was found to be a giant, while the other four were classified as MW-foreground dwarfs. We also report the presence of two other MW-contaminants classified as K V and G V stars, all confirmed in the l.o.s. velocity determination step. We have therefore excluded these MW-foreground stars from the counts in Table 2 and the following analysis. Fig. 4 shows the position of the classified stars on the colour-magnitude diagram, colour-coded according to their , visually confirming the general goodness of the spectral classification.
3.1.2 O and Be stars
@@
On the other hand, O stars are typically not uniformly distributed in space. They tend to form and evolve in OB associations. Our observations cover some of the OB associations identified by Garcia et al. (2009). These associations are low density groups of young stars with significant internal extinction, making the identification of O types more challenging. In Fig. 5 we show the spatial distribution of the OB associations reported by Garcia et al. (2009), together with that of our identified OB stars with K and also the Be stars. We see that they generally tend to be found where the OB associations are. Their spatial distribution also generally follows that of the ionised shells shown in Fig. 2, to which they are probably physically associated (Borissova et al. 2004; Garcia et al. 2009). At this stage, these evidences at least confirm the goodness of our initial classification of many of them as hot stars.
@@ 3.1.3 C/M fraction
3.2 Velocity determination
3.2.1 Velocity determination with SPEXXY
@@
The determination of l.o.s. velocities was carried out using the spexxy101010https://github.com/thusser/spexxy code (v. 2.5, Husser 2012). The routine performs a full spectral fitting to the observed spectra using interpolated spectral templates generated from the PHOENIX library of high-resolution synthetic spectra111111http://phoenix.astro.physik.uni-goettingen.de/ (Husser et al. 2013). The library covers a wide wavelength range, going from 500 Å to 5.5 m, and stellar parameters: , , and . However, the PHOENIX models do not treat radiative transfer in the atmospheres of massive stars affected by strong stellar winds, so the library is of little use in the recovery of stellar parameters for K. We refer to Appendix B.1 for a detailed explanation of the spectral fitting steps and to B.2 for a comparison of its performance with the ULySS results presented above.
@@
We ran spexxy on all targets having S/N (in at least one of the S/N indicators) and clean photometry (see Sect. 2.2). We also excluded all those targets that in the spectral classification step (see Sect. 3.1.1) were identified as contaminated by ionised gas, background galaxies, and Be stars (which we analysed separately). These steps reduced our sample to 1745 input sources. We were thus able to recover their l.o.s. velocity, along with estimates of their spectral parameters and associated errors. We note that only for of stars spexxy failed to perform a successful fitting, mostly due to their low S/N.
@@
The goodness of the recovered parameters was quantified in a series of tests, whose details are given in Appendix B.1. Here we report the essentials. We verified by performing several tests on mock spectra that the l.o.s. velocity errors are well estimated down to S/NCaT for red giant stars (S/N550 for MS stars), and the velocity values are recovered without significant offsets. On the other hand, the recovery of the spectral parameters was more limited, with in particular being generally well constrained for red giants, but underestimated for the hottest MS stars due to the grid limitation of our spectral models. We emphasise again that our aim was to use spexxy mainly for the determination of l.o.s. velocities.
@@
Since our main goal is to analyse not only the global kinematic properties of IC 1613, but also how they change as a function of stellar types, we decided to keep the sample obtained during the spectral classification step. From this sample we further removed those targets with a velocity error km s-1, which meant excluding of sources with the least reliable velocity values (mostly low S/N MS stars). We also excluded the six targets previously classified as MW contaminants, confirming that they are foreground sources based on their l.o.s. velocities, which were consistent with 0 km s-1. Performing a cross-correlation with the third data release of the Gaia catalogue (Gaia Collaboration et al. 2023), we further confirmed their foreground nature based on their proper motion values, significantly different from 0 km s-1 in this case. With these conditions applied, our sample reduced to 727 sources to which we subsequently added the 24 Be stars whose velocity determination we describe in the next section. The distribution of l.o.s. velocity measurements for our sample is shown in Fig. 6, with histograms divided by pointing and stellar type. We note that the median velocity error was of km s-1, while the median S/NC resulted around 10.
3.2.2 Velocity determination for carbon and Be stars
@@
The l.o.s. velocities of the C stars identified by spectral classification (see Sect. 3.1.1) were calculated using spexxy as for the main sample. However, the code was not always able to match the complex spectral features of these stars. Therefore, we double-checked the spexxy velocity determination with ULySS using stellar templates from the X-Shooter library of carbon stars (Gonneau et al. 2016). We found a general agreement within the errors, confirming the goodness of the spexxy velocities. The median velocity error of the 14 C stars analysed (6 in both F1 and F2, and 2 in F3) was km s-1 for a median S/N.
@@
On the other hand, the velocity determination for the Be stars was done separately, since spexxy failed in this task because such stars are not represented by the PHOENIX model atmospheres. We thus performed a cross-correlation with custom-made template spectra using the function correlation of the python package specutils, part of the Astropy project (Astropy Collaboration et al. 2022). Templates were generated as ad-hoc proxies using the spectra of two B-type stars (classified as B9III and B3III) from the MIUSCAT library (Vazdekis et al. 2012), to which we added two Gaussian profiles reproducing the strength and width of the Hα and Hβ emission lines as detected in our sample. Since the spectra of our Be-star candidates in general did not show any split line profiles (except for two objects), the two-component Gaussian approximation seemed to be a good enough approach for the purpose of l.o.s. velocity measurements. We cross-correlated the observed spectra with the templates and assigned them the average of the derived velocities. The errors were instead assigned by a Monte Carlo process. For each template we generated mock spectra at different S/N (calculated around 5500 Å) values . For each S/N, we cross-correlated 250 mock spectra with the corresponding noise-free template and considered the median absolute deviation of the velocity distribution as the velocity error representative of that bin. The error distribution as a function of S/N was fitted by an exponential profile for each template and we used the mean profile to assign velocity errors to the observed spectra according to their S/N550. For the 24 Be stars analysed (14, 6 and 4 in fields F1, F2 and F3, respectively), we recovered a median km s-1 for a median S/N.
@@ 4 Internal kinematics
| Model | k | Bayes factor | ||
| (km s-1) | (km s-1) | (km s-1 arcmin-1) | ln() | |
| Linear rotation (HI) | 4.5 | |||
| Linear rotation (optical) | 3.5 | |||
@@
We investigate the basic internal kinematic properties of the stellar component of IC 1613, such as its systemic l.o.s. velocity, its velocity dispersion and the possible presence of velocity gradients, indicating rotation of the stellar component131313On the angular scales covered by the MUSE sample, we find the perspective rotation due to the relative motion between the galaxy and the Sun to play a negligible role with up to km s-1 (following Eq. 6 of van de Ven et al. 2006 and assuming the systemic proper motion from Battaglia et al. 2022). Velocity gradients induced by tidal disturbances from the MW or M31 are also highly unlikely..
@@
To proceed with the kinematic analysis, we first need to identify the possible contaminants left in our sample. We had already removed many of them during the spectral classification process, with the rest expected to be faint foreground stars. We used then the method outlined in Taibi et al. (2020), which allows a Bayesian kinematic analysis while assigning membership probabilities , assumed to be Gaussian, to the individual targets. This approach is based on the expectation maximisation technique presented in Walker et al. (2009).
@@
The values depend on the l.o.s. velocities of the individual targets, but also on the prior information based on their radial distances from the galaxy’s centre and on the expected l.o.s. velocity distribution of the possible contaminants. The spatial prior takes into account that the probability of membership is higher, the closer to the galaxy’s centre. In this case, we simply required that our targets should follow a monotonically decreasing radial density profile applying an isotonic regression model, with no assumptions about its functional form141414https://scikit-learn.org/stable/modules/isotonic.html.
@@
We carry out a Bayesian analysis to explore and compare different kinematic models, considering one where the internal kinematics of IC 1613’s stellar component can be purely described by random motions, and another model which also contains a rotational term. For the rotational term we used the form: , where is the angular distance from the galaxy’s centre (i.e. , where () are the tangent plane’s coordinates), is the position angle (P. A.) of the j-target, and is the P. A. of the kinematic major axis (both P. A. measured from North to East). The rotational velocity term is assumed to linearly increase with radius, that is .
@@
We note that, unless otherwise stated, for the stellar component we quote the observed rotational velocity and not its intrinsic value , as the latter requires knowledge of the inclination angle i of the angular momentum vector with respect to the observer, that is . The stellar component of LG dwarf galaxies is not in a thin disc, therefore its inclination cannot be determined without knowledge of the intrinsic 3D shape. In those systems that contain a clearly rotating HI component, as IC 1613, one can at least assume the inclination determined from the neutral gas component, under the assumptions that stars and gas share the same angular momentum vector. This is what we will do when quoting values for , or when deriving estimates for the circular velocity, using preferentially the value quoted in Table 1. In Sect. 6.2, we discuss the possibility that IC 1613 may be seen close to face-on (Read et al. 2016), and suggest a new value based on the combined kinematic properties of the different stellar tracers we analysed.
4.1 Kinematic properties of the main sample: detection of a rotation signal